OpenStudy (peperwqio):

Traveling at 55 miles per hour, how many minutes, rounded to the nearest whole number, does it take to drive 310 miles from Boston to Syracuse?

OpenStudy (peperwqio):

Correct Answer gets a medal

jimthompson5910 (jim_thompson5910):

Use the formula D = r*t where D = distance r = rate (in other words speed) t = time

jimthompson5910 (jim_thompson5910):

D = 310 miles r = 55 miles per hour t = unknown

jimthompson5910 (jim_thompson5910):

let me know what you get

OpenStudy (peperwqio):

Uh i have the answer already and you didn't use the correct method to solve the problem so the medal doesn't go up for grabs and the question is closed.